# Projective imaging of high-energy nuclei through coherent exclusive vector meson production in electron-ion collisions

### v3

@@ III Results
In this section, we focus on demonstrating that one is able to alleviate the limited $|t|$ resolution and largely restore the diffractive pattern for coherent VM production by measuring the projective $|t|$ distribution using a wedge cut. The statistical separation of coherent and incoherent production through decay patterns is less intricate and will be further explored in future work.
@@
To mimic the detector resolution, we smear the coherent $|t|$ distribution from its analytic form, i.e., black solid curve in Fig. 2, with three constant resolution values, 25, 50 and 100 MeV/$c$, independent of $|t|$. This is done by convoluting Eq. 1 with a Gaussian distribution to smear the $q_{x}$ component. Resulting $|t|$ distributions are shown in Fig. 2, where one can see the designed detector resolution (open circles) lies between 25 to 50 MeV/$c$ of resolution. For the following results, the 25 MeV/$c$ resolution is used.
@@
The smeared $|F(t=q^{2})|^{2}$ distribution as a function of $q_{x}$ and $q_{y}$ is shown in Fig. 5. Compared to Fig. 4 where no detector resolution is included, the distribution becomes more elliptical along the $q_{x}$ axis, a direct consequence of the electron momentum resolution.
@@
Figure 5: Distribution of the form factor from Eq. 1 as a function of ($q_{x}$, $q_{y}$), with 25 MeV/$c$ resolution added to the $q_{x}$ axis. A wedge cut of $\theta_{\rm max}=\pi/12$ from the $q_{y}$ axis is indicated by arrows.
@@
The projected $|t|$ distribution ($|F_{\hat{n}}(t)|^{2}$) is obtained by applying a wedge cut of $\theta_{\rm max}=\pi/12$, indicated by black arrows in Fig. 5, and shown in Fig. 6 as the dashed magenta curve. Compared to the case of no wedge cut or equivalently $\theta_{\rm max}=\pi/2$ (blue long-dashed line in Fig. 6), where the diffractive pattern is largely washed out, one can see a clear structure of peaks and valleys in the projected $|t|$ distribution for a wedge cut of $\theta_{\rm max}=\pi/12$. Such a significant improvement in resolving the diffractive pattern is critical for utilizing the $|t|$ measurement for imaging the gluon spatial distribution inside nuclei. The price one pays with the wedge cut is the loss of statistics. For example, about 83% of coherent VM events are thrown away with the $\theta_{\rm max}=\pi/12$ wedge cut for the 25 MeV/$c$ resolution. The optimal wedge selection should be determined experimentally based on the detector resolution and available statistics. For completeness, the analytical distribution of the form factor is also in Fig. 6 as the black solid curve. As expected, the diffractive patten is more distinct compared to the case with detector resolution and a wedge cut.
@@
Figure 6: Projected $|t|$ distribution with a wedge cut of $\theta_{\rm max}=\pi/12$ for coherent VM production. It is compared to the theoretical expectation of the full $|t|$ distribution as well as that with detector resolution but no wedge cut (i.e., $\theta_{\rm max}=\pi/2$). Both curves have been normalized by a factor of $A/\int|F_{\hat{n}}(t)|^{2}$ where $A=197$ for gold and $|F_{\hat{n}}(t)|^{2}$ is determine by the angle $\theta_{\rm max}$.
IV Summary and Outlook
@@
In summary, we developed a unique approach to overcome the challenges of incoherent background and limited detector resolution in utilizing coherent VM production for imaging the gluon spatial distribution in the nucleus at the future EIC. Specifically, the incoherent background can be separated from coherent VM production in collisions of transversely polarized electron beams by exploiting the angular distribution of decay daughters from VM. Moreover, the projected $|t|$ distribution perpendicular to the electron scattering plane can be used to avoid the effect of limited precision in measuring the outgoing electron’s momentum with the current EIC detector design. By applying a wedge cut in the transverse $|t|$ plane, the theoretically-predicted diffractive pattern for coherent VM production is largely restored, from which the gluon distribution in the nucleus and its dynamics can be inferred.
In future work, we will implement this approach into the simulation of the ePIC detector, the baseline detector for the EIC. We will determine the best wedge cut based on the projected EIC luminosity and realistic detector resolution, including its $|t|$ dependence, and study the residual effect from energy and angular resolutions of outgoing electrons. An unfolding procedure will be explored to account for the detector resolution in measuring the projected $|t|$ distribution. This work will open the door for the mission-critical measurement of exclusive VM production in studying gluon distributions, and enhance the scientific output of the EIC project.
Acknowledgements.

### v4

@@ 3 Results
In this section, we apply the proposed projective technique and demonstrate that one is indeed able to alleviate the limited $|t|$ resolution and largely restore the diffractive pattern for coherent VM production by measuring the projective $|t|$ distribution in a specified phase space. The statistical separation of coherent and incoherent production through decay patterns is less intricate and will be further explored in future work.
@@
To simulate detector resolution, we smear the coherent $|t|$ distribution from its analytic form, i.e., black solid curve in Fig. 2, with a $|t|$-independent resolution of 25 MeV/$c$ as inspired by the EIC detector proposal [33]. This is done by convoluting Eq. 1 with a Gaussian distribution to smear the $q_{x}$ component, and the result is shown as red dotted curve in Fig. 2.
@@
The smeared $|F(t)|^{2}$ distribution as a function of $q_{x}$ and $q_{y}$ is shown in Fig. 5. Compared to Fig. 4 where no detector resolution is included, the distribution becomes more elliptical along the $q_{x}$ axis, a direct consequence of the electron momentum resolution.
@@
Figure 5: Distribution of the form factor from Eq. 1 as a function of ($q_{x}$, $q_{y}$), with 25 MeV/$c$ resolution added to the $q_{x}$ axis. A wedge cut of $\theta_{\rm max}=\pi/12$ from the $q_{y}$ axis is indicated by arrow.
@@
The projected $|t|$ distribution ($|F_{\hat{n}}(t)|^{2}$) is obtained by applying a selection of $\theta_{\rm max}=\pi/12$, indicated by the black arrow in Fig. 5, and shown in Fig. 6 as the blue dashed curve. Compared to the case without wedge cut or equivalently $\theta_{\rm max}=\pi/2$ (red dotted curve in Fig. 6), where the diffractive pattern is largely washed out, one can see a clear structure of peaks and valleys in the projected $|t|$ distribution for a wedge selection of $\theta_{\rm max}=\pi/12$. This significant enhancement in resolving the diffractive pattern is essential for using the $|t|$ distribution to image the gluon spatial distribution inside nuclei. However, it is worth noting that a loss of statistics occurs when employing a selective phase space as seen in Fig. 6. For example, about 83% of coherent VM events are excluded with the $\theta_{\rm max}=\pi/12$ wedge cut for the 25 MeV/$c$ resolution. The optimal wedge selection should be determined experimentally based on the detector resolution and available statistics. For completeness, the analytical distribution of the form factor is also in Fig. 6 as the black solid curve. As expected, the diffractive patten is more distinct compared to the case with detector resolution and without a phase-space selection in the specific $|t|$ decomposition.
@@
Figure 6: Projected $|t|$ distribution with a phase space constraint of $\theta_{\rm max}=\pi/12$ for coherent VM production. It is compared to the theoretical expectation of the full $|t|$ distribution as well as that with detector resolution but without the selection of phase space (i.e., $\theta_{\rm max}=\pi/2$).
4 Summary and Outlook
@@
In summary, we developed a unique approach to overcome the challenges of overwhelming incoherent background and limited detector resolution in utilizing coherent VM production for imaging the gluon spatial distribution inside nuclei. Specifically, the incoherent background can be separated from coherent VM production in collisions of transversely polarized electron beams by exploiting the angular distribution of decay daughters from VM. Moreover, the projected $|t|$ distribution perpendicular to the electron scattering plane can be used to avoid the effect of limited precision in measuring the outgoing electron’s momentum. This is achieved by applying a selective phase space constraint in the transverse $|t|$ plane. Consequently, the theoretically predicted diffractive pattern for coherent VM production is largely restored, enabling a high-priority program at the EIC and other facilities aimed at measuring the gluon distribution in the nucleus and its dynamics.
In future work, we will implement this approach into the simulation of the ePIC detector, the baseline detector for the EIC. We will determine the phase-space selection based on the projected EIC luminosity and realistic angular resolution of outgoing electrons. An unfolding procedure will be employed to correct for the detector resolution in the measured projected $|t|$ distribution. The proposed technique can also be applied to other electron-hadron experiments, such as those in the Continuous Electron Beam Accelerator Facility at JLab [29, 30], EicC in China [4] and LHeC at CERN [5, 6], to fully exploit the power of imaging gluon distributions with coherent VM production.